# Rep3D: Re-parameterize Large 3D Kernels with Low-Rank Receptive Modeling for Medical Imaging

### v1

@@ Appendix B Datasets Details
Challenge AMOS CT AMOS MR MSD Pancreas MSD Hepatic Vessels KiTS Imaging Modality Multi-Contrast CT Multi-Contrast MRI Venous CT Arterial CT Anatomical Region Abdomen Pancreas Liver Kidney Sample Size 361 33 282 303 300 Anatomical Label Spleen, Left & Right Kidney, Gall Bladder, Pancreas, Tumor Hepatic Vessels, Tumor Kidney, Tumor Esophagus, Liver, Stomach, Aorta, Inferior Vena Cava (IVC) Pancreas, Left & Right Adrenal Gland (AG), Duodenum Bladder (CT only), Prostate/Uterus (CT only) Data Splits 1-Fold (Internal) 5-Fold Cross-Validation Train: 160 / Validation: 20 / Test: 20 Train: 22 / Validation: 4/ Test: 7 Train: 225 / Validation: 27 / Testing: 30 Training: 242, Validation: 30 / Testing: 31 Training: 240, Validation: 30 / Testing: 30 5-Fold Ensembling N/A N/A X X
Appendix C Network Architecture
We adopt a 3D encoder-decoder architecture from both 3D UX-Net [18] and SwinUNETR [10] as the backbone of Rep3D. Instead of using encoder block with feed forward layer, we simply using a plain convolutional design with depthwise separable convolutions in parallel with LRBM. The encoder consists of 4 hierarchical stages with increasing feature dimensions and depthwise convolutions of large kernel size ($21\times 21\times 21$), followed by a symmetric decoder for volumetric segmentation. The encoder includes:
@@ Appendix C Network Architecture
Appendix D Validation Experiments on Variable Branch Learning rate
@@
Appendix E Ablation Study on Network Depth for LRBM

### v2

@@ A.3 Datasets Details
Challenge AMOS CT AMOS MR MSD Pancreas MSD Hepatic Vessels KiTS Imaging Modality Multi-Contrast CT Multi-Contrast MRI Venous CT Arterial CT Anatomical Region Abdomen Pancreas Liver Kidney Sample Size 200 33 282 303 300 Anatomical Label Spleen, Left & Right Kidney, Gall Bladder, Pancreas, Tumor Hepatic Vessels, Tumor Kidney, Tumor Esophagus, Liver, Stomach, Aorta, Inferior Vena Cava (IVC) Pancreas, Left & Right Adrenal Gland (AG), Duodenum Bladder (CT only), Prostate/Uterus (CT only) Data Splits 1-Fold (Internal) 5-Fold Cross-Validation Train: 160 / Validation: 20 / Test: 20 Train: 22 / Validation: 4/ Test: 7 Train: 225 / Validation: 27 / Testing: 30 Training: 242, Validation: 30 / Testing: 31 Training: 240, Validation: 30 / Testing: 30 5-Fold Ensembling N/A N/A X X
A.4 Network Architecture
We adopt a 3D encoder-decoder architecture from both 3D UX-Net (Lee et al., 2022) and SwinUNETR (Hatamizadeh et al., 2022a) as the backbone of Rep3D. Instead of using encoder block with feed forward layer, we simply using a plain convolutional design with depthwise separable convolutions in parallel with LRBM. The encoder consists of 4 hierarchical stages with increasing feature dimensions and depthwise convolutions of large kernel size ($21\times 21\times 21$), followed by a symmetric decoder for volumetric segmentation. The encoder includes:
@@ A.4 Network Architecture
A.5 Training Efficiency Comparison
To further validate our claims around improved convergence behavior and training efficiency, we conducted an additional ablation study focusing on runtime and performance dynamics across different configurations of Rep3D on the AMOS dataset. While our primary goal is to improve segmentation accuracy and spatial convergence, it is equally important that such gains are achieved with minimal training overhead. In this study, we compare three architectural variants:
  1. (a)
    Vanilla Rep: Rep3D with parallel convolutional branches, but without any spatial prior modulation.
  2. (b)
    Fixed Prior: Rep3D with a non-learnable reciprocal distance mask acting as a fixed spatial prior.
  3. (c)
    Full LRBM: Rep3D with a learnable low-rank bias module (LRBM), modulating the spatial prior adaptively via a generator network.
All models were trained under identical conditions: using a single NVIDIA A100 GPU, batch size of 2, and AdamW optimizer. We report the validation Dice scores at key training checkpoints (10k, 20k, 40k, and 60k iterations), as well as the total training time to convergence.
Table 7: Training efficiency and convergence comparison of different Rep3D variants on AMOS dataset.
Method Time (hrs) 10k Iter 20k Iter 40k Iter 60k Iter
Vanilla Rep 17.3 0.853 0.868 0.886 0.892
Fixed Prior 15.5 0.864 0.875 0.892 0.902
Full LRBM 17.5 0.871 0.885 0.897 0.910
As observed in Table 7, the full LRBM variant consistently achieves the best segmentation accuracy at every checkpoint, demonstrating accelerated convergence. The introduction of the learnable low-rank generator yields a modest increase in training time (+0.2+0.2 hours compared to Vanilla Rep), but this is substantially outweighed by the observed performance gains. The Fixed Prior variant also performs competitively, suggesting the benefit of incorporating even a static spatial prior.
A.6 Validation Experiments on Variable Branch Learning rate
@@
Appendix B Ablation Study on Network Depth for LRBM
@@
B.1 Additional Comparisons with nnU-Net Variants: ResEnc nnU-Net, STU-Net, MedNeXt
we conducted further comparisons with recent state-of-the-art (SOTA) 3D medical image segmentation architectures, including ResEnc nnU-Net (Isensee et al., 2021), STU-Net-H (Huang et al., 2023), and MedNeXt (Roy et al., 2023). These models have demonstrated strong performance across various benchmarks and provide important context for positioning Rep3D among contemporary architectures. All baseline methods were trained under their official recommended training schedules (typically 1000 epochs), using the same computational resources and training splits for fair comparison.
Table 10: Average Dice scores across five segmentation benchmarks under full training schedules.
Method AMOS CT AMOS MRI KiTS Pancreas Hepatic
nnU-Net (1000 epochs) 0.887 0.847 0.706 0.703 0.660
ResEnc nnU-Net 0.892 0.850 0.711 0.706 0.661
STU-Net-H 0.900 0.848 0.707 0.712 0.648
MedNeXt 0.897 0.856 0.720 0.713 0.663
Rep3D (Ours) 0.910 0.864 0.736 0.723 0.674
These comparisons further validate the strong and consistent performance of Rep3D across multiple challenging 3D segmentation benchmarks. Even under extensive training schedules (1000 epochs), Rep3D outperforms the SOTA alternatives, demonstrating its robustness and generalizability.